XRCF 2ND FLOOR SHIFT REPORT Starting Date: 20 January 1997 Shift: C File Ref: /home/xrcfdude/shiftreports/970120.C RS/RE: WAP Operator: NM Software: Shift Scientist: MS Project Scientist: SO ==================== Plan: ---- Collect Defocus and 3-D Measurements with Penning Sources. Measurement Summary: ------------------- 109658 D-LXH-3D-20.012 3-D Scan, Penning 95eV LETG HSI 109659-60 D-LXH-3D-20.010 3-D Scan, Penning 95eV LETG HSI 109661 D-LXH-3D-20.004 3-D Scan, Penning 95eV LETG HSI 109662 D-LXH-dF-7.006 3-D Scan, Penning 95eV LETG HSI 109664 D-LXH-dF-7.005 3-D Scan, Penning 95eV LETG HSI 109666 D-LXH-dF-7.007A 3-D Scan, Penning 95eV - no grating HSI 109677 D-LXH-dF-7.004 3-D Scan, Penning 95eV LETG HSI 109678 D-LXH-3D-30.001 3-D Scan, Penning 95eV LETG HSI 109679 D-LXH-3D-20.007 3-D Scan, Penning 95eV LETG HSI Results: ------- Supplementary Tests: ------------------- Notes: ----- Tube in place at beginning of shift lasted six hours; replaced tube at 07:00 in under an hour. the loc file for D-LXH-3D-30.001 was modified to add fourth point with both dy and dz < 0; we think that that the original locations were set up erroneously. Problems: -------- No photons seen in fpc detectors from 109663 - 109676; setupdet was run before starting 109677, and good spectra were then seen in them. No apparent problem with voltages, gains, etc. The origin of this problem is still a total mystery, but may indicate that blindly skipping setupdet for a large number of similar detector config measurments may not be a good idea.